MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2016/12/10
OBSERVER: Alex Li
CONDITION: Windy & Clody
COMMENT: The vacuum value was frozen at 5.9e-06 from 11:30 pm, and I recovered the vacuum reading back to the normal at 12:35 am by switching off and on the vacuum sensor. I also switched off and on the vacuum sensor again at 1:35 am because the values between the sensor and the reading on the computer seemed not matched.
Compressor gas pressure: 1.75 MPa
Expected Observation Time: START 23:04 END 04:16


Go to the Seeing Log

Go to the Observation Log

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB-000
LMC13712-1371213900
SMC-000
TRANSIT-000
ALERT-000
GRB-000
TOTAL---13900 (00h15m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
00000000000


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
010000000001
Field13Field14Field15Field16
0100


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000


TRANSIT FIELD STATISTICS

Field1Field2Field3
000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
23:04waitingwindy & cloudy
00:00waitingwindy & cloudy
01:00waitingwindy & cloudy
02:00observingcloudy, then clear for half an hour
03:00waitingcloudy
04:00waitingcloudy
04:16waitingcloudy


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day